OpenStudy (anonymous):

How do I find the antiderivative of dx / (sqrt 1-4x^2) ?

OpenStudy (anonymous):

Use a substitution, 2x=sin(u), say.

OpenStudy (anonymous):

Then\[2dx=\cos u du \]and so\[\int\limits_{}{}\frac{dx}{\sqrt{1-4x^2}}=\frac{1}{2}\int\limits_{}{}\frac{\cos u}{\sqrt{1-\sin^2 u}} du=\frac{1}{2}\int\limits_{}{} du=\frac{u}{2}+c\]

OpenStudy (anonymous):

Then \[u=\sin^{-1} \frac{x}{2}\]

OpenStudy (anonymous):

So the integral is\[\frac{1}{2}\sin^{-1} \frac{x}{2}+ c\]

OpenStudy (anonymous):

sorry, \[\sin^{-1} 2x\] I mean

OpenStudy (anonymous):

\[\int\limits_{}{}\frac{dx}{\sqrt{1-4x^2}}=\frac{1}{2}\sin^{-1} 2x +c\]

OpenStudy (anonymous):

or the other easy way is to seprate deduction like:\[1\div \sqrt{1-4*x^2}=(A/\sqrt{1-2*x})+(B/\sqrt{1+2x})\] which A=1/sqrt{2}=B

OpenStudy (anonymous):

ijadi, your method's nice

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

thanks!